5. Compute the electrostatic force experienced by -4.33 x 10-18 C from the following electricfield.

(a) E = 5 465 N/C

(b) 9.5 x 1015 N/C




1
Expert's answer
2022-02-28T11:51:55-0500

The electrostatic force

"F=qE"

(a)

"F=4.33*10^{-18}*5465=2.37*10^{-14}\:\rm N"

(b)

"F=4.33*10^{-18}*9.5*10^{15}=0.041\:\rm N"